A library crew can shelve books in eight hours. another crew can do the job in six hours. If the first crew member begins and is joined by the second crew one hour later, how many hours will it take for the job to be completed?
Answer by lwsshak3(11628)

A library crew can shelve books in eight hours. another crew can do the job in six hours. If the first crew member begins and is joined by the second crew one hour later, how many hours will it take for the job to be completed?
1/8= hourly work rate of first crew
1/6= hourly work rate of second crew
let x=hours worked by first crew
x-1 = hours worked by second crew
(1/8)x+(1/6)(x-1)=100% of the job=1
LCD=48
6x+8x-8=48
14x=56
x=4
ans: it will take 4 hours to complete the job